### v1

@@ Abstract
In object goal navigation, agents navigate towards objects identified by category labels using visual and spatial information. Previously, solely network-based methods typically rely on historical data for object affinities estimation, lacking adaptability to new environments and unseen targets. Simultaneously, employing Large Language Models (LLMs) for navigation as either planners or agents, though offering a broad knowledge base, is cost-inefficient and lacks targeted historical experience. Addressing these challenges, we present the LLM-enhanced Object Affinities Transfer (LOAT) framework, integrating LLM-derived object semantics with network-based approaches to leverage experiential object affinities, thus improving adaptability in unfamiliar settings. LOAT employs a dual-module strategy: a generalized affinities module for accessing LLMs’ vast knowledge and an experiential affinities module for applying learned object semantic relationships, complemented by a dynamic fusion module harmonizing these information sources based on temporal context. The resulting scores activate semantic maps before feeding into downstream policies, enhancing navigation systems with context-aware inputs. Our evaluations in AI2-THOR and Habitat simulators demonstrate improvements in both navigation success rates and efficiency, validating the LOAT’s efficacy in integrating LLM insights for improved object goal navigation.
Keywords:
Object Goal Navigation Commonsense Reasoning Generalization
1 Introduction
Object goal navigation is a significant area of study that focuses on enabling intelligent agents to navigate to a specified target object within a scene given category names. The objective is not only to successfully navigate to the object but also to minimize the navigation path. To make the searching process successful and more efficient, researchers have explored how to leverage object semantic relationships in household environment to assist in information filtering for more efficient navigation.
Figure 1: The Inspiration for LOAT. By extracting generalized object affinities from LLMs and experiential ones from training environments, policy with LOAT framework is able to strike a balance between generalization, computing-cost and system stability.
@@ 1 Introduction
Models relying solely on training experiences tend to overfit and underperform in new environments or with unseen objects. To cope with that, some scholars are leveraging large models for richer object contextual insights from large models to improve navigation system generalization. For instance, Zhou et. al[35] employs a pre-trained vision-language model (VLN) for open-world prompt-based positioning and a large language model to score object relevance numerically. Nonetheless, this method may face challenges like inconsistent inference scores and variable object priorities with paraphrased prompts, which could affect system stability. Empirical studies reveal that while LLMs may not deliver consistent scores across diverse prompts, they reliably identify relevant objects in a stable manner, affirming their role as a dependable source of generalized knowledge.
Explorations into directly leveraging large models (LLMs or VLMs) as navigation agents like SayNav [18] and NavGPT[34], despite promising results, the approach’s reliance on frequent and computationally intensive queries for each navigation step, limiting its practicality and application due to high time and resource demands. Moreover, while these models are adept at general reasoning, their broad knowledge base might lack the specificity required for certain navigation contexts, such as culturally specific household layouts, potentially leading to less than optimal navigational outcomes.
@@
    Integrates LLM’s commonsense reasoning with historical experiences in a dual-module framework, enhancing efficiency and generalization. Commonsense knowledge is accessible at the start of each episode or stored offline for specific tasks, offering cost-effectiveness without compromising generalization.
@@
    Showcases performance enhancements across all semantic map-based navigation policies, from metric to topological maps, within Habitat[20] and AI2-THOR[12] simulations, demonstrating its utility and applicability.
@@
    Utilizes pretrained text embeddings for target representations, decoupling target identification both from traditional one-hot encodings and environmental visuals, ensuring adaptability to unseen objects and enhancing generalization beyond specific visual features.
2 Related Work
2.1 Maps for Navigation
@@
Within SLAM research, two predominant mapping paradigms have emerged: metric based mapping and topological mapping. Metric based mapping methods [9][8][3], known for producing detailed maps by merging neighboring pixels into a grid, might face challenges related to computational complexity, particularly in extensive indoor settings [25]. These maps, while accurate, can impede efficient planning and problem-solving due to their dense representation of the environment.
Conversely, topological mapping offers a more streamlined approach by depicting the environment through interconnected nodes, representing distinct landmarks or areas [4] [33]. This representation simplifies navigation and planning by abstracting the environment into a series of relationships between key locations. However, achieving accurate and consistent topological maps remains a challenge, especially in complex environments where sensor data may not always provide clear distinctions.
@@
2.2 Navigation with Large Models Integrated
@@
Object Relevance Scoring with Scripted Strategies: This approach leverages large models to evaluate object relevance, utilizing scripted search strategies such as frontier-based searching methods[26]. For instance, ESC[35] employs LLMs for direct candidate scoring, while L3MVN[31] and Prompter[10] derive collocation probabilities from masked language models with prompts like “Something you find at [MASK] is [TARGET],” scoring candidates based on their placement in the "[MASK]" token. However, the sensitivity of these models to the nuances of language can lead to significant score variability for slight prompt modifications, affecting the consistency of object relevance assessments.
@@
3 Methods
The LLM-enhanced Object Affinities Transfer (LOAT) framework, as depicted in Fig. 2, innovatively merges LLM-derived commonsense insights with experiential navigation data to empower navigation models with object affinities augmented semantic maps. It features a generalized affinities module to harness the expansive semantic understandings of LLMs, an experiential affinities module to utilize past object relationship patterns and a dynamic fusion module to balance the combination of the two source of information given temporary contexts, significantly enhancing model adaptability and decision-making in new environments. The fused affinities are then applied for node-wise or channel-wise activation based on map categories before feeding to downstream policy.
Figure 2: Detailed LOAT Framework. To refine object affinities, LOAT uses two information sources. LLMs power the generalized affinities module with normalized binary assessment of general object relevance. The Experiential Module generates attention scores for object affinities from past training using a pretrained text embedder and key-query transformations. A temporal state embedding-based dynamic fusion module balances these information sources. Obtained scored are applied for node-wise or channel-wise activation depending on map categories before feeding downstream policy.
3.1 Experiential Affinities Module
@@
To compute the attention scores, we first transform the embeddings into queries $\mathbf{Q}$ and keys $\mathbf{K}$ using separate linear transformations for the target object and all other objects, respectively:
$\mathbf{Q}=\mathbf{W}_{q}\mathbf{e}(o_{target}),\quad\mathbf{K}=\mathbf{W}_{k}\mathbf{e}(o_{i})$ (1)
@@
$A_{E_{i}}=\frac{\exp(\mathbf{Q}\cdot\mathbf{K}^{T}/\sqrt{d_{k}})}{\sum_{j=1}^{M}\exp(\mathbf{Q}\cdot\mathbf{K}_{j}^{T}/\sqrt{d_{k}})}$ (2)
@@
3.2 Generalized Affinities Module
The generalized affinities module leverages semantic relations derived from Large Language Models (LLMs) to enhance the model’s focus on objects semantically related to a specified target. The semantic relevance of each object $o_{i}$ in the map to the target $o_{target}$ is determined by a binary value $S(o_{i},o_{target})$, indicated by an LLM. This setup ensures that all objects deemed semantically related to the target are assigned a non-zero attention weight. By employing this uniform attention mechanism, the module guarantees that the agent considers all potentially relevant objects, thus improving its adaptability and performance in unfamiliar settings.
@@ 3.2 Generalized Affinities Module
$A_{G_{i}}=\frac{S(o_{i},o_{target})}{\sum_{j=1}^{M}S(o_{j},o_{target})}$ (3)
@@
3.3 Dynamic Fusion Module
@@
The key to this adaptation is the guidance ratio $\gamma$, dynamically modulated based on the temporal context $\mathcal{H}$ and additional environmental factors $\mathcal{E}$ if available. For architectures utilizing RNNs to encode current states, $\mathcal{H}$ comprises the RNN’s hidden states, whereas for non-RNN architectures, it includes the agent’s past trajectory and relevant environmental data such as explored regions and observed layouts. $\mathcal{E}$ further enriches this context with external environmental cues, aiding in precise adjustment of $\gamma$.
This adjustment of $\gamma$ is governed by a learned function $f(o_{target},\mathcal{H},\mathcal{E})$, designed to decrease $\gamma$ when experiential patterns are reliable, such as in familiar environments, and increase it in favor of generalized affinities where semantic guidance from LLMs offers superior navigational cues, typically in novel or complex settings. The fusion of affinity scores, therefore, dynamically shifts to harness the most effective navigation strategy:
$A_{F_{i}}=\gamma\cdot A_{G_{i}}+(1-\gamma)\cdot A_{E_{i}}$ (4)
@@
3.4 Integration with Downstream Policy
@@
Map-based Policy: Let $S_{m}$ represent a metric map, with each channel assigned to pre-selected object categories, typically large and easily identifiable objects within the environment. The affinity score for each channel $c$, derived from the dynamic fusion in equation 4, is represented by $A_{F_{c}}$. The activation of channel $c$ in response to the affinity is computed as:
$Activation(c)=A_{F_{c}}\cdot S_{m}[c]$ (5)
@@
By employing these strategies, the LOAT framework significantly refines the decision-making process, enhancing the model’s adaptability and generalization capabilities across diverse environments. Through targeted activation within downstream policies, the approach leverages both learned environmental patterns and generalized semantic insights derived from large language models, culminating in a robust framework for efficient and focused navigation.
4 Experiments
4.1 Setup
4.1.1 Benchmarks:
@@
4.1.2 Evaluation Metrics:
To rigorously assess our method’s performance across benchmarks, we employ several key metrics: SR (Success Rate): Measures the percentage of episodes where the agent successfully completes the task by reaching the target object. Specifically in ALFRED, it assesses the agent’s effectiveness in both navigating to and engaging with targets to achieve task completion; PLWSR (Path Length Weighted Success Rate): Accounts for the efficiency of the agent’s path towards the target, penalizing longer paths even in successful episodes; GC (Goal Completion): A binary metric indicating whether the agent achieved the task’s objective within the given constraints; PLWGC (Path Length Weighted Goal Completion): Similar to PLWSR, this metric weights the goal completion against the optimality of the path taken; GFR (Goal Found Rate): Specifically introduced for ALFRED to capture instances where the agent locates the target object. It evaluates the method’s precision in guiding the agent to deduce the target’s location from high-level instructions, focusing on navigation success separate from interaction execution.
4.1.3 Implementation Details:
We leverage paraphrase-MiniLM-L6-v2 to compute text embeddings of category names. This BERT-based model excels in semantic similarity tasks, producing 384-dimensional dense vectors from sentences and phrases, ideal for representing object categories that may extend beyond single words.
To determine object affinities, we consult GPT-4, pre-storing semantic scores for expedited navigation. In cases of open-vocabulary navigation, this process can be executed online at each episode’s onset. This approach is feasible because the maps already have predefined and recognizable large objects to save, with the navigation target being the sole changing factor.
@@ 4.1.3 Implementation Details:
In ALFRED, we enhance the FILM[15] system with LOAT for semantic map activation, replacing MaskRCNN with a fine-tuned DINO[2] model for superior segmentation, and adjusting exploration strategy inspired by Prompter[10] for more efficient navigation. This enhanced system, termed LOAT-P, showcases an integrated approach for improved navigation and object interaction within AI2-THOR environments. LOAT-P system is contrasted with HLSM[1], which models the environment in a spatial-semantic voxel map, FILM[15], known for its spatial memory and semantic search, LGS-RPA[17] with its landmark-based guided search, EPA[14]’s neural-symbolic planning, Prompter[10]’s template-based planning, and CAPEAM[11]’s spatial and state-change integration.
@@
4.2 Results
The experiment results in ALFRED are shown in Table 1. LOAT-P achieves SOTA performance across all metrics, increasing SR by around 10%. Notably, the SR discrepancy between familiar and unfamiliar environments is markedly reduced, showcasing LOAT’s efficacy in applying generalized object affinities from LLM in novel settings. This demonstrates our system’s robust adaptability and effectiveness in diverse scenarios.
Table 1: Comparison with the state-of-the-art Methods in ALFRED.

### v2

@@ Abstract
In object goal navigation, agents navigate towards objects identified by category labels using visual and spatial information. Previously, solely network-based methods typically rely on historical data for object affinities estimation, lacking adaptability to new environments and unseen targets. Simultaneously, employing Large Language Models (LLMs) for navigation as either planners or agents, though offering a broad knowledge base, is cost-inefficient and lacks targeted historical experience. Addressing these challenges, we present the LLM-enhanced Object Affinities Transfer (LOAT) framework, integrating LLM-derived object semantics with network-based approaches to leverage experiential object affinities, thus improving adaptability in unfamiliar settings. LOAT employs a dual-module strategy: a generalized affinities module for accessing LLMs’ vast knowledge and an experiential affinities module for applying learned object semantic relationships, complemented by a dynamic fusion module harmonizing these information sources based on temporal context. The resulting scores activate semantic maps before feeding into downstream policies, enhancing navigation systems with context-aware inputs. Our evaluations conducted in the AI2-THOR and Habitat simulators indicate significant improvements in both navigation success rates and overall efficiency. Furthermore, the system performs effectively when deployed on a real robot without requiring additional training, thereby validating the efficacy of LOAT in integrating LLM insights for enhanced object-goal navigation.
Index Terms:
Object Goal Navigation, Commonsense Reasoning, Generalization
I Introduction
Object goal navigation is a significant area of study that focuses on enabling intelligent agents to navigate to a specified target object within a scene given category names. The objective is not only to successfully navigate to the object, but also to minimize the navigation path[1, 2]. To make the searching process successful and more efficient, researchers have explored how to leverage object semantic relationships in household environment to assist in information filtering for more efficient navigation.
TABLE I: Comparison of Navigation Methods
Method Traditional Navigation Methods Network Agent w LOAT
LM Agent/Planner Network Agent
Generalization
Cost-efficient
Stability
@@ I Introduction
Models relying solely on training experiences often suffer from overfitting and tend to underperform when applied to novel environments or unfamiliar objects. To cope with that, some scholars are leveraging large models for richer object contextual insights to improve navigation system generalization. For instance, Zhou et. al[7] employs a pre-trained vision-language model (VLN) for open-world prompt-based positioning and a large language model to score object relevance numerically. Nonetheless, this method may face challenges like inconsistent inference scores and variable object priorities with paraphrased prompts, which could affect system stability. Empirical studies reveal that while LLMs may not deliver consistent scores across diverse prompts, they reliably identify relevant objects in a stable manner. Therefore, we focus on object relevance rather than the specific scores provided by LLMs to ensure a more consistent and reliable source of generalized knowledge.
Explorations into directly leveraging large models (LLMs or VLMs) as navigation agents like SayNav [8] and NavGPT[9] have shown promising results. However, the approach’s reliance on frequent and computationally intensive queries for each navigation step poses significant challenges, limiting its practicality due to high time and resource demands. Moreover, while these models are adept at general reasoning, their expansive knowledge base may lack the specificity needed for certain navigation contexts—such as culturally specific household layouts—potentially resulting in suboptimal navigational performance.
@@
    Dual-Module Framework Integration: This work integrates commonsense reasoning from LLM with historical experiences, enhancing both efficiency and generalization. Commonsense knowledge is accessible at the beginning of each episode or stored offline for specific tasks, providing cost-effectiveness without compromising generalization.
@@
    Utilization of Pretrained Text Embeddings: By employing pretrained text embeddings for target representations, this approach decouples target identification from traditional one-hot encodings and environmental visuals. This ensures adaptability to unseen objects and enhances generalization beyond specific visual features.
@@
    Performance Improvements in Navigation Policies: The proposed method demonstrates significant performance enhancements across all semantic map-based navigation policies, from metric to topological maps, in simulations using Habitat[10] and AI2-THOR[11]. Furthermore, in empirical experiments, the system performs effectively when deployed on a real robot without requiring additional training, validating the efficacy of LOAT in integrating LLM insights for enhanced object-goal navigation.
II Related Work
II-A Maps for Navigation
@@
Within SLAM research, two predominant mapping paradigms have emerged: metric based mapping and topological mapping. Metric based mapping methods [13, 14, 15, 16], which generate detailed maps by aggregating neighboring pixels into a grid, are known for their precision. However, these methods may encounter challenges related to computational complexity, especially in large-scale indoor environments [17]. While these maps offer high accuracy, their dense representation of the environment can hinder efficient planning and problem-solving.
Conversely, topological mapping offers a more streamlined approach by depicting the environment through interconnected nodes, representing distinct landmarks or areas [18, 4, 19, 20]. This representation simplifies navigation and planning by abstracting the environment into a series of relationships between key locations. However, achieving accurate and consistent topological maps remains a challenge, especially in complex environments where sensor data may not consistently provide clear distinctions.
@@
II-B Navigation with Large Models Integrated
@@
Object Relevance Scoring with Scripted Strategies: This approach leverages large models to evaluate object relevance, utilizing scripted search strategies such as frontier-based searching methods[21]. For instance, ESC[7] employs LLMs for direct candidate scoring, while L3MVN[22] and Prompter[23] derive collocation probabilities from masked language models with prompts like “Something you find at [MASK] is [TARGET],” scoring candidates based on their placement in the ”[MASK]” token. However, the sensitivity of these models to subtle changes in language can result in significant variability in scores with slight prompt modifications, thereby affecting the consistency of object relevance assessments.
@@
Figure 1: Detailed LOAT Framework. In the LOAT framework, the target object and the categories of objects within the scene map are processed to determine the target location based on object affinities. Information about these categories is fed into two main modules. The Generalized Affinities Module, powered by LLMs, provides a normalized binary assessment of each object’s general relevance. Simultaneously, the Experiential Module utilizes a pretrained text embedder and key-query transformations to generate attention scores reflecting object affinities based on historical training data. These scores from both modules are then dynamically balanced by a fusion module that incorporates temporal state embeddings. Obtained scored are applied for node-wise or channel-wise activation depending on map categories before feeding downstream policy.
III Methods
Instead of employing Large Language Models (LLMs) as high-level planners, which typically necessitates complex alignments with low-level policies, our approach enables the direct utilization of language-based priors in a numerical format that is intuitively interpretable to downstream models. As illustrated in Fig. 1, the LOAT framework merges LLM-derived commonsense insights with historical object affinity data to enhance semantic maps, effectively bridging the gap between high-level semantic understanding and low level network-based policy. This is achieved by feeding target object categories and environmental data into two sub-modules: the Generalized Affinities Module, which captures broad semantic knowledge from LLMs, and the Experiential Affinities Module, which leverages learned object relationship patterns from training data. A Dynamic Fusion Module dynamically balances these sources based on temporary contexts, improving the system adaptability and decision-making capabilities in unfamiliar environments. The obtained affinity scores are then utilized for node-wise or channel-wise activation according to the map types. The enriched semantic map is then fed to downstream navigation policies for enhanced context-aware decision planning.
III-A Experiential Affinities Module
@@
To compute the attention scores, we first transform the embeddings into queries $\mathbf{Q}$ and keys $\mathbf{K}$ using learned linear transformations for the target object and all other objects respectively:
$\mathbf{Q}=\mathbf{W}_{q}\mathbf{e}(o_{target}),\quad\mathbf{K}=\mathbf{W}_{k}\mathbf{e}(o_{i}),$ (1)
@@
$A_{E_{i}}=\frac{\exp(\mathbf{Q}\cdot\mathbf{K}^{T}/\sqrt{d_{k}})}{\sum_{j=1}^{M}\exp(\mathbf{Q}\cdot\mathbf{K}_{j}^{T}/\sqrt{d_{k}})},$ (2)
@@
III-B Generalized Affinities Module
The generalized affinities module leverages semantic relations derived from Large Language Models (LLMs) to enhance the model’s focus on objects semantically related to a specified target. While an intuitive approach might be to have these large models score object affinities directly, a key challenge emerges due to the inconsistency in inference scores and varied object priorities when prompting. Paraphrased prompts can lead to unstable and inconsistent scoring, potentially undermining the system’s reliability. To address this, we shift the focus from relying on specific scores to emphasizing object relevance, ensuring a more consistent and dependable source of generalized knowledge. Therefore, we use priors from large models solely as identifiers of relevant objects rather than for exact scoring. The semantic relevance of each object $o_{i}$ in the map to the target $o_{target}$ is determined by a binary value $S(o_{i},o_{target})$, indicated by a LLM. This setup ensures that all objects deemed semantically related to the target are assigned a non-zero attention weight and remain unaffected by the object affinities from training data. By employing this uniform attention mechanism, the module guarantees that the agent considers all potentially relevant objects, thus improving its adaptability and performance in unfamiliar settings.
@@ III-B Generalized Affinities Module
$A_{G_{i}}=\frac{S(o_{i},o_{target})}{\sum_{j=1}^{M}S(o_{j},o_{target})}.$ (3)
@@
III-C Dynamic Fusion Module
@@
This adaptation is driven by the guidance ratio $\gamma$, which is dynamically modulated based on the temporal context $\mathcal{H}$ and, when available, additional environmental factors $\mathcal{E}$.
In architectures utilizing RNNs to encode the current state, $\mathcal{H}$ includes the hidden states of the RNNs. For non-RNN architectures, $\mathcal{H}$ encompasses the agent’s past trajectory and relevant environmental data, such as explored regions and observed layouts. $\mathcal{E}$ further enriches this context with external cues, aiding in the precise adjustment of $\gamma$.
The key idea involves extracting features that capture historical and environmental contexts, with processing tailored to the specific modality of each context as shown in Fig. 2. For example, if the context is a vector—such as the hidden states of RNN, which inherently represent past experiences—it is processed through a MLP. Conversely, If the context is a matrix—such as a 2D grid map representing the environment layout—it is first encoded with convolutional layers, then flattened into a vector and subsequently processed through a MLP. All available context features are concatenated and fed into the final MLP to obtain the dynamic rate $\gamma$. Building on this design, $\gamma$ is dynamically adjusted according to the context: it decreases in familiar environments where experiential patterns are consistent, and increases in novel or complex settings where LLMs provide more effective navigational insights. This dynamic fusion of affinity scores enables the system to adaptively employ the most effective navigation strategy for the given scenario:
$A_{F_{i}}=\gamma\cdot A_{G_{i}}+(1-\gamma)\cdot A_{E_{i}}.$ (4)
@@
Figure 2: Detailed Dynamic Fusion Module Architecture. The input of the dynamic fusion module could be several environmental and temporal contexts, the encoders of which are dependent on their modalities. All of the flattened features will be further concatenated together before undergoing the final MLP and outputting the guidance rate.
III-D Integration with Downstream Policy
@@
Map-based Policy: Let $S_{m}$ represent a metric map, with each channel assigned to pre-selected object categories, typically large and easily identifiable objects within the environment. The affinity score for each channel $c$, derived from the dynamic fusion in Equ. (4), is represented by $A_{F_{c}}$. The activation of channel $c$ in response to the affinity is computed as:
$Activation(c)=A_{F_{c}}\cdot S_{m}[c].$ (5)
@@
Training LOAT with Downstream Policy: The original optimization methods for downstream policies can vary, ranging from imitation or reinforcement learning to other training approaches. As long as the training process is gradient-based, LOAT can be integrated with the policy network to facilitate the training process. The training occurs in two phases, using the same dataset as the one employed for training the original policy.
In the first phase, we integrate the experimental affinity module with the base policy and train the entire network using the original loss function. In the second phase, we introduce the generalized affinity module and dynamic fusion layer, updating only the parameters of the dynamic fusion layer while keeping all other parameters frozen.
By employing these strategies, the LOAT framework significantly refines the decision-making process, enhancing the model’s adaptability and generalization capabilities across diverse environments. Through targeted activation within downstream policies, the approach leverages both learned environmental patterns and generalized semantic insights derived from large language models, culminating in a robust framework for focused navigation.
IV Experiments
IV-A Setup
IV-A1 Benchmarks
@@
IV-A2 Evaluation Metrics
To rigorously assess our method’s performance across benchmarks, we employ several key metrics: SR (Success Rate): Measures the percentage of episodes where the agent successfully completes the task by reaching the target object. Specifically in ALFRED, it assesses the agent’s effectiveness in both navigating to and engaging with targets to achieve task completion; SPL (Success weighted by Path Length): Defined as the ratio of the shortest path length to the path actually taken by the agent, weighted by whether the episode was successful, evaluating the efficiency of navigation. PLWSR (Path Length Weighted Success Rate): Accounts for the efficiency of the agent’s path towards the target, penalizing longer paths even in successful episodes; GC (Goal Completion): A binary metric indicating whether the agent achieved the task’s objective within the given constraints; PLWGC (Path Length Weighted Goal Completion): Similar to PLWSR, this metric weights the goal completion against the optimality of the path taken; GFR (Goal Found Rate): Specifically introduced for ALFRED to capture instances where the agent locates the target object. It evaluates the method’s precision in guiding the agent to deduce the target’s location from high-level instructions, focusing on navigation success separate from interaction execution.
IV-A3 Implementation Details
We leverage paraphrase-MiniLM-L6-v2 [29] to compute text embeddings of category names. This BERT-based model excels in semantic similarity tasks, producing 384-dimensional dense vectors from sentences and phrases, ideal for representing object categories that may extend beyond single words.
To determine object affinities, we consult GPT-4, pre-storing semantic scores for expedited navigation. In cases of open-vocabulary navigation, this process can be executed online at each episode’s onset. This approach is feasible because the maps already have predefined and recognizable large objects to save, with the navigation target being the sole changing factor. When dynamically calculating the guidance ratio $\gamma$, we employ two 2-layer MLPs with ReLU activation.
@@ IV-A3 Implementation Details
In ALFRED, we enhance the FILM[32] system with LOAT for semantic map activation, replacing MaskRCNN with a fine-tuned DINO[33] model for superior segmentation, and adjusting exploration strategy inspired by Prompter[23] for more efficient navigation. This enhanced system, termed LOAT-P, showcases an integrated approach for improved navigation and object interaction within AI2-THOR environments. The LOAT-P system is compared with several established methods, including HLSM [34], which employs a spatial-semantic voxel map for environment modeling; FILM [32], recognized for its spatial memory and semantic search capabilities; LGS-RPA [35], which utilizes landmark-guided search; EPA [36], featuring neural-symbolic planning; Prompter [23], which relies on template-based planning; CAPEAM [37], which integrates spatial and state-change information; and LLM-Planner [24], a training-free approach that leverages large language models for few-shot planning.
@@
IV-B Results
TABLE II: Comparison with the state-of-the-art Methods in ALFRED.